# Making Teams and Influencing Agents: Efficiently Coordinating Decision Trees for Interpretable Multi-Agent Reinforcement Learning

### v1

@@ 4.3 Validation Rollouts: UCB Policy Selection
7:     Set $m^{*}\leftarrow\operatorname{argmax}_{m}\hat{\mu}_{\ell}^{m}+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}}}$
@@
Following $M$ iterations, HYDRAVIPER produces $M$ joint policy profiles for each team. It then needs to select the best-performing policy profile (Algorithm 1, lines 13–14). HYDRAVIPER iterates through the policy profiles to estimate the team performance of each one using a set of validation rollouts. The performance metric it uses is the undiscounted mean return of the team, $\bar{R}^{m}_{\ell}=\frac{1}{T}\sum_{t=0}^{T}R_{\ell}(s_{t},\hat{\pi}^{m}_{\ell}(\mathbf{o}_{t}))$.
@@
As is the case for training, collecting validation rollouts is computationally intensive, so these rollouts also need to be efficiently allocated. However, the problem setting differs here. Our goal is not to collect a diverse set of training rollouts, but rather to identify the most performant policy profiles using as few rollouts as possible. The mean return of each policy profile is unknown a priori; it must be estimated by selecting policy profiles and performing rollouts with noisy returns. Again, a fixed allocation strategy of $K_{\mathrm{valid}}$ environment rollouts for each policy profile is wasteful. The rollouts assigned to clearly badly-performing policy profiles could be reallocated to reduce the variance in the estimated returns of promising policy profiles. This motivation aligns with that of multi-armed bandit (MAB) problems.
@@
Given a limited budget of $B_{\mathrm{valid}}$ rollouts, we represent the task of selecting the best-performing policy profile as a MAB problem. For each team $\mathcal{T}_{\ell}$, the policy profile $\hat{\pi}_{\ell}^{m}$ from each iteration $m$ is an arm, and its return is a random variable $\bar{R}_{\ell}^{m}$ with unknown mean $\mu_{\ell}^{m}$. Each rollout samples from one of these random variables, which captures the distribution of returns from environment and policy randomness. The objective is to identify the best arm $m_{\ell}^{*}=\operatorname{argmax}_{m}\mu_{\ell}^{m}$ in as few rollouts as possible, i.e. to minimise the regret with respect to the policy that selects $m_{\ell}^{*}$ for every rollout.
In this work, we use a modification of the UCB1 algorithm (Auer, Cesa-Bianchi, and Fischer 2002). This allows us to achieve logarithmic regret given a readily satisfiable assumption: that the returns $\bar{R}_{\ell}^{m}$ of the arms are bounded. Although UCB assumes that the arms are bounded in $[0,1]$, it can be modified in a manner equivalent to rescaling the rewards to remain in $[0,1]$. We rely on the general form of the Chernoff-Hoeffding bound:
Theorem 2.
Corollary 1.
This is the same bound demonstrated for UCB1, and the same holds for the lower confidence bound $\mu_{\ell}^{m}-\sqrt{\frac{c\ln B}{n_{m}}}$. Overall, this choice of $c$ leads to the same $O(\log B)$ regret bound as UCB1. However, HYDRAVIPER can also be extended to use other MAB algorithms. If the mean returns of each policy profile are assumed to be normally distributed, the UCB1-Normal algorithm (Auer, Cesa-Bianchi, and Fischer 2002) could be used; it also achieves logarithmic regret. This algorithm effectively chooses $c$ to be proportional to the arms’ sample variance; the greater the variance, the wider the confidence bound. It is also possible to obtain an offline (but biased) estimate of the sample variance by performing expert rollouts before running UCB.
@@
Given a total budget of $B_{\mathrm{valid}}$ validation rollouts, HYDRAVIPER performs them as follows. For each policy profile, it first performs $n_{\min}=\left\lceil{2\ln B_{\mathrm{valid}}}\right\rceil$ rollouts to generate initial estimates of the mean returns (Algorithm 4, lines 2–3). To allocate the remainder of the budget (lines 4–8), HYDRAVIPER follows UCB1 to select the policy profile index for the $k$th validation rollout as
@@
$m_{\ell k}^{*}=\operatorname{argmax}_{m}\left(\hat{\mu}_{\ell}^{m}(k)+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}(k)}}\right),$
@@ 4.4 Agent Clustering: Scaling Up HYDRAVIPER
When resampling the dataset, HYDRAVIPER calculates sample weights following Equation 1. This computation requires enumerating joint actions $\mathbf{a}_{\ell}$ for each team $\mathcal{T}_{\ell}$, in order to find the worst-case joint action that minimises the team-averaged $Q$-function, $\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))$. The complexity of this step scales with the size of the joint action space, and thus exponentially with the size of the team. Some mixed competitive-cooperative environments (see Section 5.1) have inherent team structure that can reduce this complexity. In cooperative environments such as TSC, HYDRAVIPER clusters the agent set into teams to improve training efficiency (Algorithm 3).
@@
Our goal is to find a clustering of the agent set into teams $\mathcal{T}_{1}\ldots\mathcal{T}_{L}$, so that HYDRAVIPER-trained DT students have performance similar to training on the full agent set, but improved scalability in that the number of actions to enumerate per team is much smaller than the full agent set: $\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\ll\prod_{i\in\{1,\ldots,N\}}\mathcal{A}_{i}$. We leverage the intuition that agents distant from each other (in terms of environmental distance, trajectory similarity, or other metrics) are unlikely to be influential on each other in most environments.
@@
How can the distance metric $d$ be defined? For an environment that has inherent team structure (such as the physical deception environment in Section 5.1), we define the graph $G$ to consist of a complete subgraph for each team:
$d_{ij}=\begin{cases}1&\textrm{if $i$ and $j$ share a team}\\ \infty&\textrm{otherwise}\end{cases}$
For traffic signal control environments, we note that the road network inherently forms a graph $G_{env}$, which can be partitioned to obtain sets of spatially proximal agents that correspond to neighbouring intersections. In this case, we simply set $G=G_{env}$, or equivalently
$d_{ij}=\begin{cases}1&\textrm{$e_{ij}\in G_{env}$}\\ \infty&\textrm{$e_{ij}\notin G_{env}$}\end{cases}$
@@ 5 Experiments
Now, we demonstrate the utility of HYDRAVIPER for interpretable MARL using experiments on various benchmark environments. In doing so, we perform a functionally grounded evaluation of interpretability (Doshi-Velez and Kim 2017), where we assess the quality of the generated DTs in terms of performance and computational efficiency. As the DTs would be used directly in lieu of NN-based policies in deployment, we consider these to be good proxy metrics of their practical applicability. More specifically, we address the following research questions:
@@ 5.1 Environments
We evaluate HYDRAVIPER on four environments: two environments in the multi-agent particle world (MPE) benchmark (Lowe et al. 2017), and two traffic signal control (TSC) environments in the RESCO benchmark (Ault and Sharon 2021). In MPE environments, agents must navigate in a 2D space to accomplish a coordinated objective, making these environments ideal for assessing coordination capabilities.
@@
In this environment, a team of two defender agents must cooperate to protect two targets from an adversary agent. One of the two targets is the “goal” for the adversary; this is not known to the adversary, which can only observe the positions of the targets and defenders. We train the two defender agents against an NN adversary.
In TSC environments, each agent controls a single intersection by selecting different signal phases; each phase permits vehicles from a subset of lanes to pass through the intersection. Both environments are based on real-world road corridors reproduced in the traffic simulator SUMO (Alvarez Lopez et al. 2018). To interface with the simulator, we use the OpenAI Gym-style wrapper sumo-rl (Alegre, Bazzan, and da Silva 2021). We focus on imitating experts for all agents as a team.
@@ 5.2 Baselines and Setup
We compare HYDRAVIPER with IVIPER and MAVIPER, which represent the state of the art in interpretable multi-agent RL with DT surrogate policies. We additionally compare with expert policies — MADDPG (Lowe et al. 2017) for MPE and MPLight (Chen et al. 2020) for TSC — and an additional baseline, imitation DT. Imitation DT does not use students to collect rollouts, nor does it perform dataset resampling; it collects the same number of training rollouts as the other algorithms and trains DTs on the collected dataset. As imitation DT performs worse than the other algorithms by a wide margin, we do not include it in Table 1 or Figure 2 but show its performance in Appendix B.
For MPE environments, we use a horizon of 25 timesteps per episode, and we trained MADDPG for 60 000 episodes as the expert for the DT students to imitate. For TSC environments, we use a horizon of 125 timesteps per episode (each timestep represents 20 seconds of simulation time), and we trained MPLight for 500 episodes as the expert. All imitation learning algorithms were run for 100 iterations to produce DTs with a maximum depth of 4. IVIPER and MAVIPER ran $K_{\mathrm{train}}=K_{\mathrm{valid}}=50$ training and validation rollouts per iteration for MPE (including for the initial iteration where rollouts are collected by the experts), and 10 rollouts per iteration for TSC. Imitation DT ran the same numbers of training rollouts. We set these to equalise the number of environment interactions per iteration.
We repeated all experiments 10 times with different random seeds, and we report the mean and 95% confidence interval of the reward over 10 rollouts performed with the final student policy profiles generated from these runs. Most experiments were run in parallel on a server with 56 2.75GHz AMD EPYC 7453 processors and 252 GiB of RAM. For these experiments, we report the number of rollouts collected, not runtimes, as the time per rollout is roughly constant. However, we also report runtimes for the execution of IVIPER, MAVIPER, and HYDRAVIPER on all four environments. For these experiments, we use the kernprof profiler (v4.1.3) to run them in sequence, with no other concurrent processes running except system routines. These experiments were run on another server with 8 4.2GHz Intel i7-7700K processors and 62 GiB of RAM.
@@ 5.3 Results
Environment Expert IVIPER MAVIPER HYDRAVIPER HYDRAVIPER (LB)
@@
Runtime (s) N/A 75 709.6 $\pm$ 441.3 52 316.9 $\pm$ 6 280.2 11 203.9 $\pm$ 67.2 6 490.0 $\pm$ 98.0
Table 1: Performance and runtimes (means and 95% confidence intervals) for HYDRAVIPER and baselines. All algorithms are given the same environment interaction budget, except for low-budget (LB) HYDRAVIPER (which uses 20% of the rollouts for MPE, 10% of the rollouts for TSC). HYDRAVIPER achieves or exceeds the performance of MAVIPER using a fraction of the runtime, and still performs well in the low-budget setting. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Appendix C shows runtimes for individual algorithm steps.
@@
HYDRAVIPER achieves strong, coordinated performance in a computationally efficient manner. First, we assess HYDRAVIPER’s performance as we vary it between two environment interaction budget levels, high (5 000 training/5 000 validation rollouts for MPE, 1 000 training/1 000 validation rollouts for TSC) and low (500 training/1 500 validation rollouts for MPE, 100 training/100 validation rollouts for TSC). As shown in Table 1, HYDRAVIPER students perform better than or comparable to students trained by the most performant DT baseline (MAVIPER for MPE, IVIPER for TSC) on all environments at both budget levels. HYDRAVIPER’s performance is also better than or comparable to the NN experts for all environments except cooperative navigation, in which all DT-based algorithms cannot achieve expert-level performance. In physical deception, although neither MAVIPER nor HYDRAVIPER substantially outperforms IVIPER given the considerable stochasticity in the environment, HYDRAVIPER achieves a level of performance much closer to MAVIPER, while its training time is an order of magnitude less than MAVIPER.
In the TSC environments, HYDRAVIPER is the best-performing algorithm at both high and low interaction budget levels. Notably, HYDRAVIPER at the high budget level substantially outperforms the expert on the Ingolstadt corridor. By contrast, MAVIPER fails to coordinate the intersection agents and is overall the worst-performing algorithm. HYDRAVIPER more than halves the runtime from both IVIPER and MAVIPER on both TSC environments.
Figure 2: Performance of HYDRAVIPER and baselines as the number of rollouts decreases. Top shows decreasing training rollouts; bottom shows decreasing validation rollouts. HYDRAVIPER’s performance stays consistent as the number of rollouts decreases. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Bars show 95% confidence intervals based on 10 randomly-seeded runs. Full results are shown in Table 2 in Appendix B.
@@
As the environment interaction budget decreases, HYDRAVIPER still outperforms baselines. Now, we investigate the ability of HYDRAVIPER to adapt to increasing budget constraints for environment interaction, as would be imposed by users who wish to quickly iterate on DT policy training. As shown in Figure 2, HYDRAVIPER’s performance in all four environments does not substantially change as the training and validation rollout budgets are reduced individually. Furthermore, in all four environments, HYDRAVIPER achieves performance on par with or better than MAVIPER at all budget levels. Therefore, HYDRAVIPER is able to maintain a Pareto frontier in the tradeoff between performance and computational efficiency.
In cooperative navigation, the performance of both HYDRAVIPER and MAVIPER remain similar as the training and validation budgets are reduced individually. However, when both budgets are reduced simultaneously (shown in Figure 5), the performance of HYDRAVIPER but not MAVIPER remains essentially unchanged. In physical deception, HYDRAVIPER still performs well even as its training budget is reduced by a factor of 10, whereas MAVIPER performs substantially worse. Furthermore, the 95% confidence intervals of HYDRAVIPER’s rewards are smaller than MAVIPER at all validation budget levels. Thus, HYDRAVIPER is able to identify performant policy profiles more consistently than MAVIPER.
@@
Even when the agent set is decomposed through clustering, HYDRAVIPER maintains its performance. Finally, we evaluate the effect of agent clustering on the performance of HYDRAVIPER. For physical deception, HYDRAVIPER by default runs with the adversary and the two defenders on opposing teams, and Table 1 and Figure 2 show that it achieves good performance. For the Ingolstadt corridor environment in the low-budget setting (100 training/100 validation rollouts), we evaluate the effect of directly clustering the agent set into two teams based on the road network graph $G_{env}$. As shown in Figure 3, the graph partitioning method introduced in Section 4.4 allows HYDRAVIPER to retain its performance even when the size of the agent set is approximately halved for each team. This reduces the size of the joint action set for the largest team from 1 944 to 72, and the total runtime of HYDRAVIPER by 20%. This is a notable improvement given the relatively small size of the dataset in the low-budget setting.
Our graph partitioning method also outperforms two baselines. First, its mean reward and reward variance are both much smaller than a random agent clustering method, in which agents are randomly assigned to one of two teams. Second, graph partitioning delivers a small performance improvement over the contiguous clustering method, which randomises between all possible cuts of the road network graph into two connected components. The strong performance of this baseline suggests that incorporating environmental structure is important to a high-quality clustering. Unlike the two baselines, however, our graph partitioning method always yields a clustering that is evenly distributed in size. The resulting clustering also corresponds to the cut under the contiguous method with the best performance.
Clustering Method Runtime (s) Worst-Performing
@@
HYDRAVIPER LB 6 490.0 $\pm$ 98.0 s [[1,2,3,4,5,6,7]]
+ random 5 130.7 $\pm$ 104.2 s [[1,2,6],[3,4,5,7]]
+ contiguous 5 135.6 $\pm$ 99.0 s [[1,2,3,4,5],[6,7]]
+ graph-partition 5 181.8 $\pm$ 149.3 s [[1,2,3],[4,5,6,7]]
Figure 3: (L) Performance of HYDRAVIPER on the Ingolstadt corridor (IC) under different agent clustering methods at the low budget level. (R) Runtimes and resulting clusterings (worst-performing clustering across 10 different random seeds shown) of HYDRAVIPER under these methods. For the clusterings, the intersection agent numbers follow those shown in Figure 1. Note that graph-partition achieves the best performance and runtime.

### v2

@@ 4.3 Validation Rollouts: UCB Policy Selection
7:  Set $m^{*}\leftarrow\operatorname*{argmax}_{m}\hat{\mu}_{\ell}^{m}+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}}}$
@@
Following $M$ iterations, HYDRAVIPER produces $M$ joint policy profiles for each team. It then must select the best-performing policy profile (Algorithm 1, lines 13–14). HYDRAVIPER iterates through the policy profiles to estimate the team performance of each using a set of validation rollouts. The performance metric it uses is the undiscounted mean return of the team, $\bar{R}^{m}_{\ell}=\frac{1}{T}\sum_{t=0}^{T}R_{\ell}(s_{t},\hat{\pi}^{m}_{\ell}(\mathbf{o}_{t}))$.
@@
As is the case for training, collecting validation rollouts is computationally intensive, so these rollouts also need to be efficiently allocated. However, the problem setting differs here. Our goal is not to collect a diverse set of training rollouts, but rather to identify the most performant policy profiles using as few rollouts as possible. The mean return of each policy profile is unknown a priori; it must be estimated by selecting policy profiles and performing rollouts with noisy returns. Again, a fixed allocation strategy of $K_{\mathrm{valid}}$ environment rollouts for each policy profile is wasteful. The rollouts assigned to clearly poorly performing policy profiles could be reallocated to reduce the variance in the estimated returns of promising policy profiles. This motivation aligns with that of multi-armed bandit (MAB) problems.
@@
Given a limited budget of $B_{\mathrm{valid}}$ rollouts, we represent the task of selecting the best-performing policy profile as a MAB problem. For each team $\mathcal{T}_{\ell}$, the policy profile $\hat{\pi}_{\ell}^{m}$ from each iteration $m$ is an arm, and its return is a random variable $\bar{R}_{\ell}^{m}$ with unknown mean $\mu_{\ell}^{m}$. Each rollout samples from one such random variable, which captures the distribution of returns from environment and policy randomness. The objective is to identify the best arm $m_{\ell}^{*}=\operatorname*{argmax}_{m}\mu_{\ell}^{m}$ in as few rollouts as possible, i.e. to minimise the regret with respect to the policy that selects $m_{\ell}^{*}$ for every rollout.
@@
In this work, we use a modification of the UCB1 algorithm (Auer, Cesa-Bianchi, and Fischer 2002). This allows us to achieve logarithmic regret given a readily satisfiable assumption: that the returns $\bar{R}_{\ell}^{m}$ of the arms are bounded (see Appendix B). Given a total budget of $B_{\mathrm{valid}}$ validation rollouts, HYDRAVIPER performs them as follows. For each policy profile, it first performs $n_{\min}=\left\lceil{2\ln B_{\mathrm{valid}}}\right\rceil$ rollouts to generate initial estimates of the mean returns (Algorithm 4, lines 2–3). To allocate the remainder of the budget (lines 4–8), HYDRAVIPER follows UCB1 to select the policy profile index for the $k$th validation rollout as
@@
$m_{\ell k}^{*}=\operatorname*{argmax}_{m}\left(\hat{\mu}_{\ell}^{m}(k)+\sqrt{\frac{c\ln B_{\mathrm{valid}}}{n_{m}(k)}}\right),$
@@ 4.4 Agent Clustering: Scaling Up HYDRAVIPER
When resampling the dataset, HYDRAVIPER calculates sample weights following Equation 1. This computation requires enumerating joint actions $\mathbf{a}_{\ell}$ for each team $\mathcal{T}_{\ell}$, in order to find the worst-case joint action that minimises the team’s mean $Q$-function, $\min_{\mathbf{a}_{\ell}}\bar{Q}^{\pi_{\ell}^{*}}(\mathbf{o}_{k},\mathbf{a}_{\ell},\pi^{*}_{-\ell}(\mathbf{o}_{-\ell k}))$. The complexity of this step scales with the size of the joint action space and thus exponentially with the size of the team. Some mixed competitive-cooperative environments (see Section 5.1) have an inherent team structure that can reduce this complexity. In cooperative environments such as TSC, HYDRAVIPER clusters the agent set into teams to improve training efficiency (Algorithm 3).
@@
Our goal is to find a clustering of the agent set into teams $\mathcal{T}_{1}\ldots\mathcal{T}_{L}$ so that HYDRAVIPER-trained DT students have performance similar to those trained on the full agent set, but improved scalability in that the number of actions to enumerate per team is much smaller than the full agent set: $\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\ll\prod_{i\in\{1,\ldots,N\}}\mathcal{A}_{i}$. We leverage the intuition that agents distant from each other (in terms of environmental distance, trajectory similarity, or other metrics) are unlikely to be influential on each other in most environments.
@@
How can the distance metric $d$ be defined? For an environment that has an inherent team structure (such as the physical deception environment in Section 5.1), we define the graph $G$ as a complete subgraph for each team. For traffic signal control environments, we note that the road network inherently forms a graph $G_{env}$, which can be partitioned to obtain sets of spatially proximal agents that correspond to neighbouring intersections. In this case, $G=G_{env}$.
Environment-agnostic distance metrics can also be designed. We consider a measure of proximity that aligns with the VIPER framework: the influence of agents on each other’s $Q$-values. Recall from Section 4.1 that HYDRAVIPER measures the importance of samples by the mean $Q$-function $\bar{Q}^{\pi_{\ell}^{*}}$. In a cooperative setting, if one agent’s actions have a significant impact on another agent’s $Q$-values, then the joint actions of these agents are likely to have a significant impact on the overall agent set’s mean $Q$-function. Including both agents on the same team would allow HYDRAVIPER to capture the effects of these joint actions. We define $G$ using the distance function
$d(i,j)$ $=\mathbb{E}_{(\mathbf{o},\mathbf{a})\in\mathcal{D}_{\textrm{dist}}}\frac{2}{\delta_{ij}+\delta_{ji}},$
$\delta_{ij}$ $=Q^{\pi_{j}^{*}}(\mathbf{o},\pi^{*}(\mathbf{o}))-\min_{a_{i}}Q^{\pi_{j}^{*}}(\mathbf{o},a_{i},\pi^{*}_{-i}(\mathbf{o}_{-i})),$
where $\delta_{ij}$ is the range in agent $j$’s $Q$-values induced by agent $i$, and the distance $d_{ij}$ is the inverse of the average of $\delta_{ij}$ and $\delta_{ji}$. This symmetrisation of influence is a simplifying assumption to obtain a single weight for each edge. Alternative distance metrics could be designed to better capture agent pairs where one is significantly more influential than the other. Since METIS requires integral edge weights, we rescale $\delta_{ij}$ to percentiles between $\min_{i,j}\delta_{ij}$ and $\max_{i,j}\delta_{ij}$.
@@ 5 Experiments
Now, we demonstrate the utility of HYDRAVIPER for interpretable MARL using experiments in various benchmark environments. In doing so, we perform a functionally grounded evaluation of interpretability (Doshi-Velez and Kim 2017), where we assess the quality of the generated DTs in terms of performance and computational efficiency. As the DTs would be used directly in place of NN-based policies in deployment, we consider these to be good proxy metrics for their practical applicability. More specifically, we address the following research questions:
@@ 5.1 Environments
We evaluate HYDRAVIPER in four environments: two environments in the multi-agent particle world (MPE) benchmark (Lowe et al. 2017), and two traffic signal control (TSC) environments in the RESCO benchmark (Ault and Sharon 2021). In MPE environments, agents must navigate in a 2D space to achieve a coordinated objective, making these environments ideal for assessing coordination capabilities.
@@
In this environment, a team of two defender agents must cooperate to protect two targets from an adversary agent. One of the two targets is the “goal” of the adversary; this is not known to the adversary, which can only observe the positions of the targets and defenders. We train the two defender agents against an NN adversary.
In TSC environments, each agent controls a single intersection by selecting different signal phases; each phase allows vehicles from a subset of lanes to pass through the intersection. Both environments are based on real-world road corridors reproduced in the traffic simulator SUMO (Alvarez Lopez et al. 2018). To interface with the simulator, we use the OpenAI Gym-style wrapper sumo-rl (Alegre, Bazzan, and da Silva 2021). We focus on imitating experts for all agents as a team.
@@ 5.2 Baselines and Setup
We compare HYDRAVIPER with IVIPER and MAVIPER, which represent the state of the art in interpretable multi-agent RL with DT surrogate policies. In addition, we compare with expert policies — MADDPG (Lowe et al. 2017) for MPE and MPLight (Chen et al. 2020) for TSC — and an additional baseline, imitation DT. Imitation DT does not use students to collect rollouts, nor does it perform dataset resampling; it collects the same number of training rollouts as the other algorithms and trains DTs on the collected dataset. As imitation DT performs worse than the other algorithms by a wide margin, we do not include it in Table 1 or Figure 3 but show its performance in Appendix C.
For MPE environments, we use a horizon of 25 timesteps per episode, and we trained MADDPG for 60 000 episodes as the expert for the DT students to imitate. For TSC environments, we use a horizon of 125 timesteps per episode (each timestep represents 20 seconds of simulation time), and we trained MPLight for 500 episodes as the expert. All imitation learning algorithms were run for 100 iterations to produce DTs with a maximum depth of 4. IVIPER and MAVIPER ran $K_{\mathrm{train}}=K_{\mathrm{valid}}=50$ training and validation rollouts per iteration for MPE (including for the initial iteration where rollouts are collected by the experts), and 10 rollouts per iteration for TSC. Imitation DT ran the same number of training rollouts. We set these to equalise the number of environment interactions per iteration.
We repeated all experiments 10 times with different random seeds, and we report the mean and 95% confidence interval of the reward over 10 rollouts performed with the final student policy profiles generated from these runs. Most experiments were run in parallel on a server with 56 2.75GHz AMD EPYC 7453 processors and 252 GiB of RAM. For these experiments, we report the number of rollouts collected, not runtimes, as the rollout time is roughly constant. However, we also report runtimes for the execution of IVIPER, MAVIPER, and HYDRAVIPER on all four environments. For these experiments, we use the kernprof profiler (v4.1.3) to run them in sequence, with no other concurrent processes running except system routines. These experiments were run on another server with 8 4.2GHz Intel i7-7700K processors and 62 GiB of RAM.
@@ 5.3 Results
Environment Expert IVIPER MAVIPER HYDRAVIPER HYDRAVIPER LB
@@
Runtime (s) N/A 75 927.5 $\pm$ 203.3 58 676.4 $\pm$ 1 392.3 11 263.8 $\pm$ 55.8 6 462.4 $\pm$ 30.9
Table 1: Performance and runtimes (means and 95% confidence intervals) for HYDRAVIPER and baselines. All algorithms are given the same environment interaction budget, except for low budget (LB) HYDRAVIPER (which uses 20% of the rollouts for MPE, 10% of the rollouts for TSC). HYDRAVIPER achieves or exceeds the performance of MAVIPER using a fraction of the runtime, and still performs well in the low budget setting. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Appendix D shows runtimes for individual algorithm steps.
@@
HYDRAVIPER achieves strong, coordinated performance in a computationally efficient manner. First, we assess HYDRAVIPER’s performance as we vary it between two environment interaction budget levels, high (5 000 training/5 000 validation rollouts for MPE, 1 000 training/1 000 validation rollouts for TSC) and low (500 training/1 500 validation rollouts for MPE, 100 training/100 validation rollouts for TSC). As shown in Table 1, HYDRAVIPER students perform better than or comparable to students trained by the most performant DT baseline (MAVIPER for MPE, IVIPER for TSC) in all environments at both budget levels. HYDRAVIPER’s performance is also better than or comparable to the NN experts for all environments except cooperative navigation, in which all DT-based algorithms cannot achieve expert-level performance. In physical deception, although neither MAVIPER nor HYDRAVIPER substantially outperforms IVIPER given the considerable stochasticity in the environment, HYDRAVIPER achieves a level of performance much closer to MAVIPER, while its training time is an order of magnitude shorter than MAVIPER.
In TSC environments, HYDRAVIPER is the best performing algorithm at both the high and low interaction budget levels. Notably, HYDRAVIPER at the high budget level substantially outperforms the expert on the Ingolstadt corridor. By contrast, MAVIPER fails to coordinate the intersection agents and is in general the worst-performing algorithm. HYDRAVIPER more than halves the runtime of both IVIPER and MAVIPER on both TSC environments.
In Figure 2, we show a decision tree generated by HYDRAVIPER for one agent in the Ingolstadt corridor at the low budget level (100 training/100 validation rollouts). Even with this limited environment interaction budget, HYDRAVIPER generates DTs that are not only performant, but also intuitively sensible. Each internal node of the DT compares the number of queueing vehicles for some turning movement to a threshold; the left branch includes all samples where there are fewer vehicles than the threshold, while the right branch includes all samples where there are more than the threshold. The root assesses traffic from the westbound side street. If it is low, the DT coordinates north-south traffic on the main road; if it is high, the DT coordinates turn traffic from the side street. Stakeholders, such as traffic engineers, can follow such a workflow to visualise, reason about, and supervise DTs generated by HYDRAVIPER.
Figure 2: Decision tree generated by HYDRAVIPER (100 training/100 validation rollouts) for the Ingolstadt corridor (IC).
Figure 3: Performance of HYDRAVIPER and baselines as the number of rollouts decreases. Top shows decreasing training rollouts; bottom shows decreasing validation rollouts. HYDRAVIPER’s performance stays consistent as the number of rollouts decreases. For physical deception, higher rewards are better; for all other environments, lower rewards are better. Bars show 95% confidence intervals based on 10 randomly-seeded runs. Full results are shown in Table 2 in Appendix C.
@@
As the environment interaction budget decreases, HYDRAVIPER still outperforms baselines. Now, we investigate the ability of HYDRAVIPER to adapt to increasing budget constraints for environment interaction, as would be imposed by users who wish to quickly iterate on DT policy training. As shown in Figure 3, HYDRAVIPER’s performance in all four environments does not change substantially as the training and validation rollout budgets are individually reduced. Furthermore, in all four environments, HYDRAVIPER achieves performance on par with or better than MAVIPER at all budget levels. Therefore, HYDRAVIPER is able to maintain a Pareto frontier in the tradeoff between performance and computational efficiency.
In cooperative navigation, the performance of both HYDRAVIPER and MAVIPER remains similar as the training and validation budgets are reduced individually. However, when both budgets are reduced simultaneously (shown in Figure 5), the performance of HYDRAVIPER but not MAVIPER remains essentially unchanged. In physical deception, HYDRAVIPER still performs well even as its training budget is reduced by a factor of 10, whereas MAVIPER performs substantially worse. Furthermore, the 95% confidence intervals of HYDRAVIPER’s rewards are smaller than those of MAVIPER at all validation budget levels. Thus, HYDRAVIPER is able to identify performant policy profiles more consistently than MAVIPER.
@@
Even when the agent set is decomposed through clustering, HYDRAVIPER maintains its performance. Finally, we evaluate the effect of agent clustering on the performance of HYDRAVIPER. For the Ingolstadt corridor environment in the high budget setting (1 000 training/1 000 validation rollouts), we evaluate two strategies from Section 4.4: (1) clustering the agent set into two teams based on the road network graph $G_{env}$ (graph-metis), and (2) using pairwise $Q$ values to identify mutually impactful agents, and either $k$-means clustering (marginal-kmeans) or METIS (marginal-metis) for partitioning.
As shown in Figure 4, these clustering strategies allow HYDRAVIPER to retain its performance even when the size of the agent set is approximately halved for each team. Clustering reduces the largest team’s joint action set in size from 1 944 to 72, and the total runtime of HYDRAVIPER by up to 47%. The best-performing strategy combines pairwise $Q$-value weights with METIS for partitioning, but they all perform similarly to the unpartitioned algorithm.
These clustering methods also outperform two baselines. First, the random baseline randomly assigns each agent to one of two teams; this baseline has very high variance in performance. Second, the contiguous baseline uses a handcrafted division of the agent set into two subsets; graph-metis recovers this division automatically, but marginal-metis improves further by grouping agents that are not adjacent in the environment.
Clustering Method Runtime (s) Worst
@@
HYDRAVIPER HB 11 263.8 $\pm$ 55.8 [[1,2,3,4,5,6,7]]
+ random 6 925.5 $\pm$ 897.0 [[1,2,3,6,7],[4,5]]
+ contiguous 6 731.3 $\pm$ 214.3 [[1,2,3],[4,5,6,7]]
+ graph-metis 6 026.4 $\pm$ 39.1 [[1,2,3],[4,5,6,7]]
+ marginal-kmeans 8 538.0 $\pm$ 643.0 [[1,2,3,4,6,7],[5]]
+ marginal-metis 6 067.5 $\pm$ 57.0 [[1,2,4,7],[3,5,6]]
Figure 4: (L) Performance of HYDRAVIPER on the Ingolstadt corridor (IC) under different agent set clustering methods at the high budget (HB) level. (R) Runtimes and worst-performing clusterings across 10 different random seeds of HYDRAVIPER under these methods. The intersection agent numbers follow those shown in Figure 1.